OpenStudy (anonymous):

FG = (3 - 1)2 + (-1 - 3)2

OpenStudy (anonymous):

What are you asking for?

OpenStudy (anonymous):

how to solve it

OpenStudy (anonymous):

What is FG

OpenStudy (anonymous):

5

OpenStudy (anonymous):

3-1 = 2, -1 - 3 = -4 2^2 + -4^2 4 + 16 20

myininaya (myininaya):

-4^2 is not the same as (-4)^2 he meant (-4)^2 not -4^2

OpenStudy (anonymous):

ok thx y isnt anyone on the biology one i need help for EOCT

myininaya (myininaya):

i don't know anything about biology

OpenStudy (anonymous):

ok thx any way